OpenStudy (anonymous):

\sum_{k=1}^{10} [\sin (2k \pi/11) +i\cos (2k \pi/11) ]

OpenStudy (anonymous):

\sum_{k=1}^{10} [\sin (2k \pi/11) +i\cos (2k \pi/11) ]

OpenStudy (anonymous):

\[\sum_{k=1}^{10} [\sin (2k \pi/11) +i\cos (2k \pi/11) ]\]

OpenStudy (anonymous):

\[\sum_{k=1}^{10} [\sin (2k \pi/11) +i\cos (2k \pi/11) ]\]

OpenStudy (anonymous):

I guess it doesn't work when you are entering your question. So do you know what \[\sum_{k=1}^{10} [\sin (2k \pi/11)\] is?

OpenStudy (anonymous):

yeah that should be zero right?

OpenStudy (anonymous):

Yeah. So what is \[\sum_{k=1}^{10} [\cos (2k \pi/11)\]

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

It is cos(2pi/11) + cos(4pi/11) + ... + cos(20pi/11) which I don't like at all. Are you sure you have the right question?

OpenStudy (anonymous):

yeah.. isn't cos(2kpi) =1?? where k is any constant

OpenStudy (anonymous):

Apparently it is -10 I plugged it into MAthematica to be honest. So 0 is the real part and -10 is the imaginary part. I'm not aware of any easy way of solving the summation of cos(2kpi/11)

OpenStudy (anonymous):

Uh quick question. Is the /11 outside or inside the cosine?

OpenStudy (anonymous):

you know what my book says the correct answer is -i. i dont know how the hell that shows up!

OpenStudy (anonymous):

inside. \[\sum_{k=1}^{10} [\sin (2k \pi/11) +i\cos (2k \pi/11) ]\]

OpenStudy (anonymous):

Weird, mathematica is telling me otherwise. Sorry.